OpenStudy (anonymous):

True or false, A negative exponent will turn positive when the reciprocal is found?

OpenStudy (anonymous):

Trueski. \[\left(a^{-1}\right) \Rightarrow \left(\frac{1}{a^{-1}}=(a^{-1})^{-1}=a\right)\]